OpenStudy (anonymous):

Two buses leave a station at the same time and travel in opposite directions. One bus travels 9 mi faster than the other. If the two buses are 702 mi apart after 6 hours, what is the rate of each bus?

OpenStudy (anonymous):

51.75 mph slower bus? 60.75 faster bus? How would someone go a decimal of a speed?

OpenStudy (nurali):

Distance each bus travels is given as d=tr where - d is distance bus travels - t is a time bus travels - r is a speed of the bus : Distance first bus travels for 4 hours is d1=6r. Distance other bus travels for 4 hours is d2=6(r+9). : Total distance between them is 702 miles and since they go in opposite directions this distance is sum of distances each bus travels. : d1+d2=702 6r+6(r+9)=702 6r+6r+54=702 12r+54=702 12r=702-54 12r=648 r=648/12 r=54 mi, this is a speed of the first bus : r+9=63 mi, this is a speed of the second bus

OpenStudy (anonymous):

I was adding instead of subtracting I think.. Here is what I did 6 hrs 9 (slower) 6 hrs +9 mph (faster) = 81 6 hrs both went 702 mi 6s + 6s + 81 = 702 12s = 702 12s= 702 - 81 = 621 12s = 621 12/621 = 51.75 mph (slower) 51.75+ 9mph = 60.75 mph faster